the electric potential 5.54 m from a charge is -3680 V what is the value of the charge


1
Expert's answer
2021-02-18T18:53:53-0500

By the definition of the electric potential, we have:


"V=k\dfrac{Q}{r},""Q=\dfrac{Vr}{k}=\dfrac{-3680\ V\cdot 5.54\ m}{9\cdot10^9\ \dfrac{Nm^2}{C^2}}=-2.26\cdot10^{-6}\ C=-2.26\ \mu C."